OpenStudy (anonymous):

In a certain class election, Rhonda received 6 out of every 8 votes cast.How many votes did she get if 320 students voted? A. 240 votes B. 200 votes C. 220 votes D. 330 votes E. None of the choices are correct Sorry but I am not good in word problems :I

OpenStudy (anonymous):

6/8 = x/320 cross multiply 8x = (320)(6) 8x = 1920 x = 1920/8 x = 240 she got 240 votes